# Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond

### v2

@@ 3.2 Skeleton Contrastive Learning
Our pipeline for contrastive learning follows MoCo V2 [26]. For a positive pair $(z,z^{\prime})$, the model optimizes the InfoNCE objective defined in Eq. (2) along with a distillation objective in Eq. (4) as regularization:
@@
Meanwhile, Mixing augmentations are adopted to construct inter-transformed views. Specifically, given two skeleton sequences $s_{1},s_{2}$, we obtain the mixed skeleton data $s_{inter}$, e.g., by Mixup. Subsequently, we can obtain the embeddings corresponding to the mixed data by $z_{inter}=h_{q}\circ f_{q}(s_{inter})$. The optimized objective is for $(z_{inter},z_{inter}^{\prime})$, where $z_{inter}^{\prime}=(1-\lambda)(h_{k}\circ f_{k}(s_{1}))+\lambda(h_{k}\circ f_{k}(s_{2}))$.
2) Temporal Asymmetric Clip Contrastive Learning. The above designs focus on the sequence-level consistency modeling, i.e., the augmented data often contains sufficient temporal motion information. However, it ignores the short-term motion representation learning of model, which is necessitated for downstream tasks with dense prediction such as action detection. Meanwhile, the consistency learning for short clips, which can be viewed as a challenging temporal augmentation, can also benefit the whole representation quality learned by model. To this end, we present the clip-level skeleton contrastive learning design in this part.
To sample a motion sub-sequence (clip), we can apply the Temporal Crop transformation, and the obtained clip is much shorter ($10\%\sim 40\%$) than the original sequence. However, the clip only contains partial motion information and directly aligning the semantic embeddings of two clips is difficult for model, leading to the unstable training. Therefore, we introduce an asymmetric design, which only feeds the short clip into query branch, while the key branch takes the normal original sequence as an anchor. Meanwhile, we keep the two augmented data possessing the same spatial views to decouple the temporal variation. Specifically, we first apply the intra-augmentations to a sample $s$ to generate the $s_{ach}$. Then we further utilize the Temporal Crop to sample a continuous clip $s_{clip}$ from $s_{ach}$. It means that, $s_{ach}$ is subsuming $s_{clip}$ in temporal dimension while sharing exactly the same spatial transformations, to provide a precise and feasible target for clip-level consistency learning. Similarly, the model takes the positive pair $(z_{clip},z_{ach})$ for optimization, where $z_{clip}=h_{q}\circ f_{q}(s_{clip})$. In implementation, we just mask $z_{ach}=z^{\prime}$ for efficiency.
@@ 3.3 Masked Skeleton Prediction
$\mathcal{L}_{Mask}=\mathbb{E}\left(||(s-dec\circ f_{q}(s_{mask}))\odot(\mathbf{1}-M)||_{2}\right),$ (6)
@@ 3.4 On the Connection of Contrastive Learning and Masked Prediction
As discussed in the work [98], simply combining these two tasks can be sub-optimal due to the inherent gap of their feature modeling mechanisms [57, 23]. To this end, we explore the potential synergy for exploitation between them.
@@
(((( $z_{mask},z^{\prime}),(z_{predict},z^{\prime})\}.$
@@ 3.5 The Whole Training Strategy
$\mathcal{L}_{Info}^{query}$ and $\mathcal{L}_{KD}^{query}$ are the respective terms for the specific $query$ view summarized in Eq. (7). During training, the following objective is applied to the whole model:
@@
2) Post-Distillation Refinement. To further improve the representation quality, we introduce a post-distillation strategy as an optional refinement process. After obtaining a good prior feature space by Eq. (8), we directly remove the one-hot label constraint in InfoNCE objective $\mathcal{L}_{Info}^{all}$, and only apply the soft label, i.e., the distillation loss $\mathcal{L}_{KD}^{all}$, which assigns the attraction weights adaptively according to the similarity. This can be seen as a more explicit feature clustering process to obtain a more compact representation space by alleviating the false negative problem.
@@ 4 Benchmark Skeleton SSL Methods
Table 1: Summary of popular skeleton datasets. * denotes the collected by Marker-MoCap methods.
Dataset Instance Class Sub. View Joint
CMU* [13] 2235 45 144 - 31
SBU [92] 300 8 7 - 15
NW-UCLA [76] 1,494 10 10 3 20
UWA3D [58] 1,075 30 10 5 15
SYSU [28] 480 12 40 1 20
NTU 60 [62] 56,880 60 40 80 25
PKUMMD [43] 21,545 51 66 3 25
Kinetics [32] $\sim$260k 400 - - 18
TSU [14] 16,115 31 18 - 17
NTU 120 [44] 114,480 120 106 155 25
UAV [39] 22,476 155 119 17
BABEL* [56] 48,978 120 - - 25
4.1 Datasets
We first summarize the popular datasets for skeleton action understanding in Table 1. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
1) NTU RGB+D 60 Dataset (NTU 60) [62] is the most popular skeleton dataset. There are 56,578 videos with 25 joints for a human, captured by three Microsoft Kinect v2 cameras. The skeleton sequences are divided into 60 action categories, performed by 40 volunteers. Two evaluation protocols are recommended: a) Cross-Subject (xsub): the data for training are collected from 20 subjects, while the other 20 subjects are for testing. b) Cross-View (xview): the training set consists of front and two side views of the action performers, while testing set includes the left and right 45 degree views.
2) NTU RGB+D 120 Dataset (NTU 120) [44] is an extension to NTU 60 dataset. There are 114,480 videos collected with 120 action categories, performed by 106 subjects. Meanwhile, 32 collection setups with respect to the location and background are used to build the dataset. Two recommended protocols are presented: a) Cross-Subject (xsub): the data for training are collected from 53 subjects, while the testing data are from the other 53 subjects. b) Cross-Setup (xset): the training data uses even setup IDs, while testing data are odd setup IDs.
3) PKU Multi-Modality Dataset (PKUMMD) [43] is another large-scale benchmark with available skeleton data. Two subsets, Part I and Part II, are provided. PKUMMD Part I contains 1,076 long video sequences, with 20 action labels per video on average, and $\sim$20,000 instances are included in 51 action categories after trimming. Part II contains 2000 short video sequences with approximately seven instances each, focusing on the short-margin action detection task. It is more challenging due to the data noise and view variation.
@@ 4.2 Model Backbones
Besides, some SSL works turn to other model backbones. Convolutional neural networks (CNNs) are utilized to process the skeleton sequence as a pseudo-2D image. Meanwhile, skeletons can also be treated as point clouds, and hence the Dynamic Graph CNN (DGCNN) [79] is also explored as the feature extractor.
@@ 4.3 Downstream Tasks
3) Action Detection is a task that detects the start and end time of actions in an untrimmed skeleton sequence as well as its corresponding action label. It can also be referred to as the Temporal Action Localization or Action Segmentation. Following the previous works [49, 43], the Average Precision (mAP) at different temporal Intersection over Union (tIoU) thresholds between the predicted and the ground truth intervals is utilized as the metric.
@@ 4.4 Implementations Details
For the proposed method, we follow the implementation and experiment settings of previous works [98, 52]. For the existing methods, we give priority to the reported results in their original paper. Meanwhile, we reproduce some results to keep the fairness of the comparison and implementation. Details can be found in Supp.
4.5 Benchmark on Action Recognition Task
We benchmark different SSL methods on NTU-60 and NTU-120 benchmarks using the widely adopted action recognition task in Table LABEL:tab:main_comp. More comprehensive comparison on additional datasets can be found in the supplementary material. The feature dimension of the last layer before the classifier is provided, which often correlates with the upper capacity of the model. For the evaluation of other tasks, the community still lacks unified benchmarks. To this end, we reproduce some popular works with the same settings as our work to obtain comparable results in Sec. 4.6.
@@ 4.5 Benchmark on Action Recognition Task
1) Comparison across Different Pretext Tasks. As shown in Fig. 5, context-based and generative learning pretext tasks are mainly studied in the earlier works first. However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design. Recently, MAMP [51] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, demonstrating promising results for Transformer-based generative learning.
@@
Contrastive learning has always been a popular topic triggered by the success in the image field, e.g., SimCLR and MoCo. The model can learn a separable high-level representation space, making it dominant in learning linear representations. These works start from the exploiting the baseline algorithms [60, 72], and significantly boost the performance by applying stronger augmentations [24, 97, 41], effective training strategies [96, 7], cross-modal knowledge [38, 52], achieving rapid improvement in recent three years.
@@
Besides, combining different pre-training paradigms [98, 42] for skeleton-based representation learning has demonstrated promising results recently. This indicates that the representations learned by different pre-training paradigms can be complementary and beneficial.
@@
2) Comparison across Different Model Backbones. Different model backbones often employ different pre-training tasks. For example, Transformer often consumes extensive computational resources, making it difficult for contrastive learning, which often requires encoding multiple data views. In contrast, it naturally fits the MSM schema which reduces computational overhead by masking tokens with large ratios. On the other hand, GRU and GCN models are relatively more efficient in training, often adopting the contrastive learning pretext task. It is also noted that the GRU models are often with larger feature dimensions than GCN models due to the smaller induced computational graph and less GPU memory occupancy.
@@
Generally, different from supervised training, GRU models have achieved state-of-the-art performance under linear evaluation as shown in Fig. 5 in SSL, and the Transformer models are also promising recently. For fully fine-tuning protocol, Transformer models are more popular due to the strong representation power. Remarkably, with self-supervised pre-training, a vanilla Transformer [52] has surpassed the GCN/Transformer with complex designs in supervised training [35, 64], demonstrating the significant role for alleviating the over-fitting problem and generalization capacity improvement of SSL.

### v3

@@ 3.2 Skeleton Contrastive Learning
Our pipeline for contrastive learning follows MoCo V2 [he2020momentum]. For a positive pair $(z,z^{\prime})$, the model optimizes the InfoNCE objective defined in Eq. (2) along with a distillation objective in Eq. (4) as regularization:
@@
The intuition behind this is to combine the one-hot hard labels in InfoNCE to explicitly cluster features while avoiding collapse, and the soft labels in the distillation objective, i.e., the similarity scores, to achieve fine-grained relational knowledge learning, which also alleviates the false negative problem as discussed in Sec. 2.2.3.(1).
@@
Figure 6: The overview of the proposed method. We integrate masked skeleton modeling and contrastive learning paradigms to learn the joint- (blue part), clip- (yellow part), and sequence-level (green part) representations. Meanwhile, a series of domain-specific prompts are introduced as training guidance.
@@
Meanwhile, Mixing augmentations are adopted to construct inter-transformed views. Specifically, given two skeleton sequences $s_{1},s_{2}$, we obtain the mixed skeleton data $s_{inter}$, e.g., by Mixup. Subsequently, we can obtain the embeddings corresponding to the mixed data by $z_{inter}=h_{q}\circ f_{q}(s_{inter})$. The optimized objective is $\mathcal{L}_{CL}(z_{inter},z_{inter}^{\prime})$ in Eq. (5), where $z_{inter}^{\prime}=(1-\lambda)(h_{k}\circ f_{k}(s_{1}))+\lambda(h_{k}\circ f_{k}(s_{2}))$.
2) Temporal Asymmetric Clip Contrastive Learning. The above designs only consider the sequence-level consistency modeling, i.e., the augmented data often contains sufficient temporal motion information. However, it ignores the short-term motion modeling, which is necessitated for downstream tasks of dense prediction such as action detection. On the other hand, the consistency learning for short-term motions, e.g., clips, which can be viewed as challenging temporal augmented views, can also benefit the whole representation quality learned by model. To this end, we present the clip-level contrastive learning design.
To sample a motion clip, we can apply the Temporal Crop transformation, and the obtained clip is much shorter ($10\%\sim 40\%$) than the original sequence. However, the clip only contains partial motion information and directly aligning the semantic embeddings of two clips is difficult for model, leading to the unstable training. Therefore, we introduce an asymmetric design, which only feeds the short clip into query branch, while the key branch takes the pre-crop original sequence as an anchor. Meanwhile, we keep the two augmented data possessing the same spatial views to decouple the temporal variation. Specifically, we first apply the intra-augmentations to a sample $s$ to generate the $s_{ach}$. Then we further utilize the Temporal Crop to sample a continuous clip $s_{clip}$ from $s_{ach}$. It means that, $s_{ach}$ is subsuming $s_{clip}$ in temporal dimension while sharing exactly the same spatial transformations, to provide a precise and feasible target for clip-level consistency learning. Similarly, the model takes the positive pair $(z_{clip},z_{ach})$ for $\mathcal{L}_{CL}$ optimization in Eq. (5), where $z_{clip}=h_{q}\circ f_{q}(s_{clip})$. In implementation, we just mask $z_{ach}=z^{\prime}$ for efficiency.
@@ 3.3 Masked Skeleton Prediction
[…]
@@ 3.4 On the Connection of Contrastive Learning and Masked Prediction
As discussed in the work [zhang2023prompted], simply combining these two tasks can be sub-optimal due to the inherent gap of their feature modeling mechanisms [qi2023contrast, gui2023survey]. To this end, we explore the potential synergy for exploitation between them.
@@
$($ $z_{mask},z^{\prime}),(z_{predict},z^{\prime})\}.$
They are fed into the encoders to optimize the $\mathcal{L}_{Info}$ and $\mathcal{L}_{KD}$ in Eq. 5.
@@
With these synergetic designs, the masked skeleton modeling provides novel positives to the contrastive learning, boosting the model generalization capacity. On the other hand, the supervision signals as gradients of contrastive learning branch also assists the masked prediction branch with the high-level semantic guidance provided by inter-sample contrast. These designs alleviate the gap of feature modeling mechanism of these two tasks, which limits the performance of previous works that directly combine them, and hence yield better representation quality.
@@ 3.5 The Whole Training Strategy
$\mathcal{L}_{Info}^{query}$ and $\mathcal{L}_{KD}^{query}$ are the respective terms for the specific queryquery view summarized in Eq. (7). During training, the following objective is applied to the whole model:
@@
2) Post-Distillation Refinement. To further improve the representation quality, we introduce a post-distillation strategy as an optional refinement process. After obtaining a good prior feature space by Eq. (8), we directly remove the one-hot label constraint in InfoNCE objective $\mathcal{L}_{Info}^{all}$, i.e.,
$\mathcal{L}=\lambda_{m}\mathcal{L}_{Mask}+\lambda_{kd}\mathcal{L}_{KD}^{all},$ (10)
where we only apply distillation term that assigns the attraction weights adaptively according to the similarity. This can be seen as a more explicit feature clustering process to obtain a more compact representation space by alleviating the false negative problem. Note this single distillation objective would not cause model collapse due to the well initialized representation space learned in previous pre-training stage.
3.6 Discussion
We summarize our method as a direct response to the challenges mentioned in previous literature review. We target at the versatile action representation learning, adopting our previous work [zhang2023prompted] as the baseline for its effectiveness to learn multi-granularity features. Building on this, we further integrate the clip-level contrastive learning with it to present a more comprehensive solution with joint, clip, and sequence level modeling.
As a response to previously discussed challenges, 1) we carefully design the spatial-temporal augmentations involving intra- and inter- data samples to present a strong contrastive learning baseline, which eases the efforts of subsequent works. 2) Further, the proposed multi-granularity representation modeling can effectively enhance the feature diversity, alleviating the dimensional collapse implicitly according to [zhang2022mask]. 3) Meanwhile, for the false negative problem, we employ a different perspective, i.e., post-training to deal with it. In the following sections, we will present the quantitative analysis on these designs.
@@ 4 Benchmark Skeleton SSL Methods
We first introduce the following datasets, which are widely used in skeleton-based SSL evaluation. The summary of more datasets for skeleton action understanding can be found in Supp.
1) NTU RGB+D 60 Dataset (NTU 60) [shahroudy2016ntu] is the most popular skeleton dataset. There are 56,578 videos with 25 joints for a human, captured by three Microsoft Kinect v2 cameras. The skeleton sequences are divided into 60 action categories, performed by 40 volunteers. Two evaluation protocols are recommended: a) Cross-Subject (xsub): the data for training are collected from 20 subjects, while the other 20 subjects are for testing. b) Cross-View (xview): the training set consists of front and two side views of the action performers, while testing set includes the left and right 45 degree views.
2) NTU RGB+D 120 Dataset (NTU 120) [liu2019ntu] is an extension to NTU 60 dataset. There are 114,480 videos collected with 120 action categories, performed by 106 subjects. Meanwhile, 32 collection setups with respect to the location and background are used to build the dataset. Two recommended protocols are presented: a) Cross-Subject (xsub): the data for training are collected from 53 subjects, while the testing data are from the other 53 subjects. b) Cross-Setup (xset): the training data uses even setup IDs, while testing data are odd setup IDs.
3) PKU Multi-Modality Dataset (PKUMMD) [liu2020pku] is another large-scale benchmark with available skeleton data. Two subsets, Part I and Part II, are provided. PKUMMD Part I contains 1,076 long video sequences, with 20 action labels per video on average, and $\sim$20,000 instances are included in 51 action categories after trimming. Part II contains 2000 short video sequences with approximately seven instances each, focusing on the short-margin action detection task. It is more challenging due to the data noise and view variation.
@@ 4.1 Model Backbones
Besides, some SSL works turn to other model backbones. Convolutional neural networks (CNNs) are utilized to process the skeleton sequence as a pseudo-2D image. Meanwhile, skeletons can also be treated as point clouds, and hence the Dynamic Graph CNN (DGCNN) [wang2019dynamic] is also explored as the feature extractor.
@@ 4.2 Downstream Tasks
3) Action Detection is a task that detects the start and end time of actions in an untrimmed skeleton sequence as well as its corresponding action label. It can also be referred to as the Temporal Action Localization or Action Segmentation. Following the previous works [luo2018graph, liu2020pku], the Average Precision (mAP) at different temporal Intersection over Union (tIoU) thresholds between the predicted and the ground truth intervals is utilized as the metric.
@@ 4.3 Implementations Details
For the proposed method, we follow the implementation and experiment settings of previous works [zhang2023prompted, mao2022cmd]. For the existing methods, we give priority to the reported results in their original paper. Meanwhile, we reproduce some results to keep the fairness of the comparison and implementation. Details can be found in Supp.
4.4 A Multi-Task Benchmark for Skeleton-based SSL
In this section, we present our multi-task benchmark for skeleton-based SSL methods, which is the first time to the best of our knowledge, to benefit the community. Considering most previous SSL works are evaluated only on action recognition, we split the benchmark into two parts for presentation, i.e., a full comparison on action recognition task and the reproduction benchmark on other tasks. First, we curate a comprehensive comparison of the widely adopted action recognition task in Table LABEL:tab:main_comp with the backbone and pretext task category mark.
On the other hand, for other different tasks, an available benchmark is provided covering action retrieval, occluded recognition, action detection, and few-shot learning tasks. We select the recent popular works with official code released, and make huge efforts to reproduce them on the same settings to present a multi-task comparison as shown in Table LABEL:tab:other_tasks.
4.4.1 Comparison on Action Recognition Task
@@
1) Across Pretext Tasks. As shown in Fig. 7, context-based and generative learning pretext tasks are mainly studied in the earlier works. However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design. Recently, MAMP [mao2023masked] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol. The subsequent works [lin2025idempotent, wu2025macdiff] also demonstrate promising results for Transformer-based MSM. However, these methods are only reported based on the joint view, which makes cross-modal knowledge ensemble in this paradigm still an open problem.
@@
Contrastive learning has always been a popular topic triggered by the success in the image field, e.g., SimCLR and MoCo. The model can learn a separable high-level representation space, making it dominant in learning linear representations. These works start from the exploiting the baseline algorithms [rao2021augmented, thoker2021skeleton], and significantly boost the performance by applying stronger augmentations [guo2022aimclr, zhang2022hiclr, lin2023actionlet], effective training strategies [zhang2022cpm, chen2023self], cross-modal knowledge [li20213d, mao2022cmd], achieving rapid improvement in recent three years.
@@
Besides, combining different pre-training paradigms [zhang2023prompted, lin2024mutual] for skeleton-based representation learning has demonstrated promising results recently. This indicates that the representations learned by different pre-training paradigms can be complementary and beneficial.
@@
2) Across Model Backbones. Different model backbones often employ different pre-training tasks. For example, Transformer often consumes extensive computational resources, making it difficult for contrastive learning, which often requires encoding multiple data views. In contrast, it naturally fits the MSM schema which reduces computational overhead by masking tokens with large ratios. On the other hand, GRU and GCN models are relatively more efficient in training, often adopting the contrastive learning pretext task. It is also noted that the GRU models are often with larger feature dimensions than GCN models due to the smaller induced computational graph and less GPU memory occupancy.
@@
Generally, different from supervised training, GRU models have achieved state-of-the-art performance under linear evaluation as shown in Fig. 7 in SSL, and the Transformer models are also promising recently. For fully fine-tuning protocol, GRU model is less effective due to the over-fitting caused by high dimension and huge parameters. Instead, Transformer models are more popular due to the powerful attention mechanism. Remarkably, with self-supervised pre-training, a vanilla Transformer [mao2023masked] has surpassed the GCN/Transformer with complex designs in supervised training [lee2023hierarchically, shi2020decoupled], demonstrating the significant role for alleviating the over-fitting problem and generalization capacity improvement of SSL.
4.4.2 Comparison on Other Downstream Tasks
Table 2: Analytics on the pre-training of different datasets for multiple downstream tasks.
Pre-training Recognition Retrieval Occluded Detection Few Shot
NTU 60 NTU 120 NTU 60 NTU 120 NTU 60 xview PKUMMD NTU 120 xset
xsub xsub xsub xsub Spatial Temporal mAP@0.1 mAP@0.5 (5,1) (5,5)
NTU 60 84.8 73.5 75.4 58.8 88.0 85.8 75.5 69.8 66.7 84.8
NTU 120 85.1 76.7 75.3 64.5 87.7 88.2 75.3 69.4 69.6 87.2
PKUMMD 77.6 67.5 65.1 52.9 78.0 79.3 70.7 65.7 58.9 78.9
NTU 120+PKUMMD 85.8 77.0 76.2 64.5 88.4 88.0 76.2 70.5 69.4 87.0
1) Experiment Setting. We introduce the implementations of each task first and more details can be found in Supp.
Skeleton-based Action Retrieval. Following previous work [su2020predict], a K-nearest neighbors (KNN) classifier (kk=1) is adopted to retrieve the nearest training sample for each testing data in the representation space. All models are first pre-trained and then evaluated on the target dataset. The precision is reported as accuracy.
Action Recognition with Occlusion. We evaluate the transfer ability of representations learned from the clean dataset to the action recognition with occluded data. We adopt the linear evaluation protocol. Following [song2019richly], we construct a synthetic occluded dataset on both spatial and temporal dimension. For spatial dimension, different body parts are randomly masked. For temporal dimension, we set a random block of frames to zeros. The testing set is constructed by the same masks across different methods, with a masking ratio of [0.3, 0.7].
Skeleton-based Action Detection. We follow the previous works [luo2018graph, liu2020pku], and evaluate the short-term motion modeling capacity by action detection task. We train the attached linear classifier to predict frame-level categories to produce the final proposal. The encoder is pre-trained on NTU 60 dataset and then we fix it, transferring the learned representations to untrimmed PKUMMD Part I dataset. The mAP of different actions is adopted as metric with different tIoUs.
Unsupervised Few-shot Learning. We evaluate the performance of SSL as few-shot learners follow the previous work [lu2022self]. Specifically, the model is first pre-trained on NTU 60 dataset. Then, a simple classifier, e.g., Support Vector Machine (SVM), is fitted on the output features by the pre-trained encoder of the support data set. Finally, the adapted classifier along with the encoder is utilized to infer the query samples. We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected. which are not seen in pre-training, as the support/query sets.
2) Comparison and Analytics. The results are summarized in Table LABEL:tab:other_tasks. We utilize the officially released model weights, keeping the experiments as fair and aligned as possible.
We first present the discussion on the perspective of model backbones. GRU models usually employ higher representation dimensions due to the smaller GPU memory consumption, which makes a more informative representation available, leading to stronger performance in retrieval task than GCNs and Transformers. GCNs model the skeleton more compactly with structured prior knowledge, resulting in a better capacity to handle over-fitting, accounting for the stronger few-shot performance for unseen categories. However, the other side of the coin is the sensitivity to occlusion, where GCN can suffer from serious representation degradation because of the distribution shift caused by occlusion. Besides, it is found the GRU-based often performs better than the GCN model in terms of action detection, which primarily focuses on the spatial structure while ignoring the fine-grained temporal dependency. In contrast, GRUs model the skeletons in a temporal-based auto-regressive manner and possess better scalability for the long-sequences and short-clips. Besides, the Transformer-based model [mao2023masked] also demonstrates promising results when employing the masked skeleton modeling pretext task, and can be expected with a better fine-tuning performance due to the powerful attention mechanism.
On the taxonomy of pretext tasks, due to the lack of promising open-sourced context-based methods, we mainly focus on contrastive and generative learning methods. First, it is found that contrastive learning methods usually perform better than reconstruction-based methods, e.g., MAMP [mao2023masked] and CRRL [qi2023contrast], in action retrieval and few-shot learning. These tasks mainly rely on the direct measure in the representation space, i.e., cosine similarity, indicating the dominant position of contrastive learning for achieving highly compact representations. Remarkably, the masked modeling method, e.g., MAMP [mao2023masked], achieves significant results in action detection task. We believe that masked modeling with Transformer architecture is very promising for handling the action detection task. By splitting the long sequences into short clips as patches for masked prediction, the learned representations can well capture fine-grained action cues and achieve notable results. However, it is not observed with as good performance under occluded recognition as expected, although it explicitly involves masked data modeling, which can be explored in future work.
4.4.3 Cross-Dataset Analytics for Different Downstream Tasks
We study the effect of the dependency relationship between different pre-training datasets on multiple downstream tasks with the proposed method. Specifically, we pre-train the model on NTU 60, NTU 120, PKUMMD, and their hybrid dataset, respectively, which is then evaluated on various downstream tasks under the transfer learning setting. To achieve a better understanding, we first introduce the relationship between different datasets in terms of the data composition. Concretely, NTU 120 extends NTU 60 on the action categories from 60 to 120 without introducing new samples of the original categories. PKUMMD is another skeleton dataset with similar collection settings, sharing a significant overlap of category distribution with NTU 60 and possessing fewer samples, which can be approximated as a smaller version of NTU 60. The quantitative results are presented in Table 2.
$\bullet$ First, as expected, reducing the scale of pre-training data generally results in performance degradation for downstream tasks, especially for the evaluation of the large-scale NTU 120 dataset. This is because the new classes in testing are unseen during pre-training, resulting in limited generalization capacity.